A bottle of medicine must be stored at a temperature between 15°C and 25°C, inclusive. What range does this correspond to on the Fahrenheit scale?
(Enter your answer using interval notation.)

Answer by stanbon(75887)

A bottle of medicine must be stored at a temperature between 15°C and 25°C, inclusive. What range does this correspond to on the Fahrenheit scale?
(Enter your answer using interval notation.)
--------
F = (9/5)C + 32
F(15) = (9/5)15 + 32 = 27+32 = 59
F(25) = (9/5)25 + 32 = 45+32 = 77
----
Ans:: [59,77]
----------
Cheers,
Stan H.
-----------